OpenStudy (anonymous):

Verify the identity: Use the fundamental , sum-difference , & double-angle identities in the steps. (sin A + cos A)² = sin 2A + 1

myininaya (myininaya):

First try expanding the (x+y)^2 part Recall (x+y)^2=x^2+2xy+y^2

myininaya (myininaya):

You will use a couple identities after and be done

OpenStudy (anonymous):

\[(sinA+cosB)^2=sin^2A+2sinAcosB+cos^2A\]

myininaya (myininaya):

Hmm... I wonder if @distinction could have done that part given the formula I gave

myininaya (myininaya):

We should probably let the askers do some of the work.